OpenStudy (anonymous):

MEDAL AND FAN WILL BE GIVEN JUST HELP! Just help me with this. Part A: The area of a square is (9x2 - 12x + 4) square units. Determine the length of each side of the square by factoring the area expression completely. Show your work. Part B: The area of a rectangle is (25x2 - 16y2) square units. Determine the dimensions of the rectangle by factoring the area expression completely. Show your work. Part C: The volume of a rectangular box is (x3 - 7x2 - 9x + 63) cubic units. Determine the dimensions of the rectangular box by factoring the volume expression completely

OpenStudy (anonymous):

@triciaal @ikram002p @DangerousJesse @ayeshaafzal221 @help_people @Willie579

OpenStudy (anonymous):

@Willie579 Off topic, but who won the Flaming Taco/Flamingo/Rhino War?

OpenStudy (willie579):

They all died. THE END.

OpenStudy (anonymous):

Yay!

OpenStudy (anonymous):

Okay, now I need help

OpenStudy (willie579):

Eh. I'll try to Google and research but I'm only in 7th grade. :/

OpenStudy (anonymous):

Nvm, it's getting answered in the Mathematics section. Thanks anyways!

OpenStudy (crabbyoldgamer):

This video will help with factoring the quadratic: https://www.youtube.com/watch?v=Y9zyVCyKno8

OpenStudy (rayaz):

\[9x^2-12x+4\] Notice how this can be factorised to give a perfect square \[(3x)^2-2(3x)(2)+(2)^2 = (3x-2)^2\]\[Area = (3x-2)^2\] In a square the area is give by squaring the length, therefore the length is \[\sqrt{(3x-2)^2}=3x-2\]

OpenStudy (anonymous):

\[Rectange\ Area= 25x^2-16x^2= (5x)^2-(4x)^2=(5x+4x)(5x-4x)\\ as \ (5x-4x)<(5x+4x)\\so\ length \ of\ rectangle=5x+4x\ and \ breadth= 5x-4x\\ Volume=x^3-7x^2-9x+63\\=x^2(x-7)-9(x-7)\\=(x^2-9)(x-7)\\=(x^2-3^2)(x-7)\\=(x+3)(x-3)(x-7)\\as\ (x+3)>(x-3)>(x-7)\\we\ can \ take \ length= (x+3), breadth= (x-3)\ height= (x-7)\ or\ as\\ given \ condition\ of\ length \ breadth\ and\ height.\]

OpenStudy (anonymous):

PART a FOLLOW RaYaZ, ALL SIDES ARE POSTIVE AS OBIVIOUS.